OpenStudy (anonymous):

find cos(270degrees-45degrees)

OpenStudy (anonymous):

Cos(270-45)=Cos(3*90-45)=sin(-45)=\[-1/\sqrt{2}\]

OpenStudy (imstuck):

Do you know your identities for situations such as these? There is an identity for cos ( A - B). You would need something like that in order to do one of these problems.